OpenStudy (anonymous):

I need to verify that these two functions are inverses. f(x) = x-7/x+3 g(x) = -3x-7/x-1

OpenStudy (anonymous):

Please help!!!!

OpenStudy (agent47):

y = (x-7)/(x+3), is that the first function?

OpenStudy (anonymous):

Yes.

OpenStudy (agent47):

To verify that the second function is the inverse, just flip the y and xs in the first one, then solve for y: x = (y-7)/(y+3) (y+3)x = (y-7) yx+3x = y-7 yx - y = -7-3x y(x-1) = -(7+3x)

OpenStudy (agent47):

y = (-7-3x)/(x-1), or y = (-3x-7)/(x-1)

OpenStudy (anonymous):

yes they both are

OpenStudy (anonymous):

The question says: Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x.

OpenStudy (anonymous):

Could you show me how to solve it like that?

OpenStudy (anonymous):

Please???